# Accelerated Sequential Flow Matching: A Bayesian Filtering Perspective

### v1

@@ Abstract
Sequential prediction from streaming observations is a fundamental problem in stochastic dynamical systems, where inherent uncertainty often leads to multiple plausible futures. While diffusion and flow-matching models are capable of modeling complex, multi-modal trajectories, their deployment in real-time streaming environments typically relies on repeated sampling from a non-informative initial distribution, incurring substantial inference latency and potential system backlogs. In this work, we introduce Sequential Flow Matching, a principled framework grounded in Bayesian filtering. By treating streaming inference as learning a probability flow that transports the predictive distribution from one time step to the next, our approach naturally aligns with the recursive structure of Bayesian belief updates. We provide theoretical justification that initializing generation from the previous posterior offers a principled warm start that can accelerate sampling compared to naïve re-sampling. Across a wide range of forecasting, decision-making and state estimation tasks, our method achieves performance competitive with full-step diffusion while requiring only one or very few sampling steps, therefore with faster sampling. It suggests that framing sequential inference via Bayesian filtering provides a new and principled perspective towards efficient real-time deployment of flow-based models.
@@ 1 Introduction
To learn the belief-to-belief probability flow, we propose a flow matching objective function that can be used to effectively finetune a pretrained flow-based model. We evaluate the proposed approach on a diverse set of streaming tasks, including forecasting, planning and control, and state estimation. Our results show that belief-update-based inference achieves a compelling trade-off between fidelity and efficiency: competitive performance to full-step diffusion with only one or a few sampling steps.
2 Preliminary
@@ 3.1 Problem Setup and Bayesian Filtering
Equations 2,3 highlight a key structural property of sequential inference: the current belief $p(x_{t}|z_{\leq t})$ can be obtained by propagating the previous belief $p(x_{t-1}|z_{\leq t-1})$ via the dynamics $p(x_{t}|x_{t-1},z_{\leq t-1})$ and a correction using new observation $z_{t}$. Although this exact formulation is often intractable due to unknown system dynamics, Bayesian filtering can be interpreted abstractly as an operator:
@@ 3.2 Sequential Inference with Flow-based Models: Existing Practices
$\frac{d}{d\tau}x_{t}(\tau)=v(x_{t}(\tau),\tau;z_{\leq t}),\quad x_{t}(0)\sim p(x_{t}|z_{\leq t}),\;\;x_{t}(1)\sim\mathcal{N}(0,I),$ (5)
@@
To deploy such a model in sequential inference, a common practice is to repeatedly apply the conditional prediction model over time $t$. This receding-horizon practice is widely adopted in time-series forecasting (Box et al., 2015) and decision-making systems such as model predictive control (Garcia et al., 1989). For flow-based models, this naïve practice requires restarting a fresh sampling process and solving ODE Eq. 5 from scratch to incorporate any new observation (Janner et al., 2022).
Consequently, this pipeline suffers from high inference latency for real-time systems: solving ODE Eq. 5 from the non-informative Gaussian demands tens or hundreds of numerical solver iterations and often fails to match the observation arrival frequency, producing potential delayed predictions. This problem becomes more severe when it comes to long-horizon prediction, as each round of estimation is more expensive to compute. Fundamentally, the restarting-sampling paradigm discards historical estimations and, by initializing each update from an independent Gaussian prior, treats temporally correlated physical processes as independent across time.
@@ 3.3 Sequential Flow Matching
From the perspective of Bayesian filtering (Eqs. 2-3), inference over future states should proceed recursively: the belief at time $t$ is obtained by refining the previous belief using new observations. This structure suggests that historical estimations should serve as prior information rather than being discarded. For example, consider a forecasting task where $x_{t}=s_{t+1:t+H}$. The prediction $x_{t-1}=s_{t:t+H-1}$ at time t1t-1 already provides a coarse estimate of future states and naturally approximates $x_{t}=s_{t:t+H}$. Ignoring such information results in “redundant” computation and unnecessary sampling overhead.
Motivated by the recursive structure, we propose to directly parameterize and learn the filtering operator in Eq. 4 via a probability flow. We introduce Sequential Flow Matching model, which learns a probability flow of predictive distribution from time t1t-1 to time $t$, conditioned on newly arrived observations $z_{<t}\rightarrow z_{\leq t}$. The key step is to replace the non-informative Gaussian distribution with the previous estimation $p(x_{t-1}|z_{\leq t-1})$ as the source distribution in flow ODE:
$\frac{d}{d\tau}x_{t}(\tau)=v(x_{t}(\tau),\tau;z_{\leq t}),\quad x_{t}(0)\sim p(x_{t}|z_{\leq t}),\;\;\bm{x_{t}(1)\sim p(x_{t-1}|z_{\leq t-1})}.$ (6)
The process of solving this ODE can be viewed as performing Bayesian filtering in Eq. 4. Figure 1 C illustrates the inference pipeline of sequential flow matching.
@@
(A) independent Gaussian coupling. Suppose $x_{t}(0)\sim p(x_{t}|z_{\leq t})$ and $x_{t}(1)\sim\mathcal{N}(0,I)$ are independently sampled. Let $p_{\text{Gaussian}}$ be the distribution of $\hat{x}_{t}(0)=x_{t}(1)-v(x_{t}(1),1;z_{\leq t})$ using one-step sampling by ODE Eq. 5. Then, $\mathcal{W}^{2}_{2}(p_{\text{Gaussian}},p)=\operatorname{Var}(x_{t}|z_{\leq t}),$ where $\mathcal{W}_{2}(p,q)$ is the 2-Wasserstein distance.
(B) temporally-correlated coupling. Suppose $(x_{t-1},x_{t})\sim p(x_{t-1}|z_{\leq t-1})p(x_{t}|x_{t-1},z_{\leq t})$, and $x_{t}(0)=x_{t},x_{t}(1)=x_{t-1}$. Let $p_{\text{Bayes}}$ be the distribution of one-step sampling by ODE Eq. 6. Then
@@ 3.4 Practical Implementation
In order to learn sequential flow matching (Eq. 6), the key ingredient is to draw samples of $x_{t-1},x_{t}$ and construct a training dataset. In this section, we discuss practical considerations for dataset construction to train a parameterized velocity $v_{\theta}$ for sequential flow matching.
@@
Distribution alignment via pretrained model rollouts. To reduce train-test mismatch, we leverage a pretrained flow-based model which better approximates $x_{t-1}$ encountered at test time. Concretely, we first pretrain a flow matching $v_{\theta_{0}}$ to learn $p_{\theta_{0}}(x_{t}|z_{\leq t})\approx p(x_{t}|z_{\leq t})$ in a standard sampling-from-Gaussian manner (Eq. 5). By letting the neural architecture of $v_{\theta_{0}}$ be the same as $v_{\theta}$, we can expect $p_{\theta_{0}}(x_{t-1}|z_{\leq t-1})$ better capture the test-time source distribution $p_{\theta}(x_{t-1}|z_{\leq t-1})$. In practice, we construct pairs $(x_{t-1},x_{t})$ where $x_{t-1}$ comes from pretrained model and $x_{t}$ comes from ground truth. This construction aligns the source distribution with the model predictive distribution encountered at test time, while avoiding compounding errors that would arise from fully synthetic trajectories. The constructed dataset is used to finetune $\theta$ initialized from $\theta_{0}$. We found this hybrid construction effectively reduce train-test mismatch and adopt it throughout our experiments.
Re-noising mechanism. While using generated data from pretrained model reduces train-test mismatch, the small deviations in model prediction at test-time can still accumulate and lead to compounding errors over multiple rounds of belief update. To mitigate the error accumulation, we adopt a re-noising mechanism that injects noise into the previous estimation of $x_{t-1}$ at both training and test time, before applying the flow. Concretely, in flow ODE Eq. 6, we replace source distribution $p(x_{t-1}|z_{\leq t-1})$ with a noisy version $\tilde{x}_{t-1}:=\alpha(\tau_{\text{renoise}})x_{t-1}+\sigma(\tau_{\text{renoise}})\cdot\mathcal{N}(0,I)$, where $\alpha(\tau_{\text{renoise}}),\sigma(\tau_{\text{renoise}})$ are the coefficients of the probability flow path at noise level (flow time) $\tau_{\text{renoise}}$. We treat $\tau_{\text{renoise}}$ as a hyperparameter that can be tuned in practice.
The re-noise level $\tau_{\text{renoise}}$ controls the trade-offs between preserving historical prediction information for sampling acceleration and accommodating uncertainty during belief refinement. If $\tau_{\text{renoise}}=1$, we have $\tilde{x}_{t-1}\sim\mathcal{N}(0,I)$, which reduces to the restarting-sampling paradigm that requires a large number of sampling steps. If $\tau_{\text{renoise}}=0$, it relies on a precise modeling of $x_{t-1}\sim p(x_{t-1}|z_{\leq t-1})$, which might induce error accumulation for long-horizon prediction. In Appendix C.4 we study the impact of $\tau_{\text{renoise}}$, and we observe that a suitable re-noise level varies with the level of environmental stochasticity.
We present the finetuning algorithm (Algorithm 1) and the inference algorithm (Algorithm 2).
@@
1:Ground-truth $x_{t}$, context $z_{\leq t}$, renoise level $\tau_{\text{renoise}}$, and fixed pretrained model $v_{\theta_{0}}$ (from Eq. 5)
2:Generate $x_{t-1}\sim p_{\theta_{0}}(x_{t-1}|z_{\leq t-1})$ using $v_{\theta_{0}}$
3:Renoise $\tilde{x}_{t-1}=\alpha(\tau_{\text{renoise}})\cdot x_{t-1}+\sigma(\tau_{\text{renoise}})\cdot\mathcal{N}(0,I)$
4:Sample interpolation time $\tau\sim\text{Uniform}(0,1)$
5:Compute $x_{t}(\tau)=\alpha(\tau)x_{t}+\sigma(\tau)\tilde{x}_{t-1}$
6:return loss $L(\theta)=\norm{v_{\theta}(x_{t}(\tau),\tau;z_{\leq t})-\dot{x}_{t}(\tau)}^{2}$
@@
1:Pretrained model $v_{\theta_{0}}$ (from Eq. 5), sequential flow matching model $v_{\theta}$ finetuned by Algorithm 1, initial observation $z_{1}$
2:initialize $x_{1}\sim p_{\theta_{0}}(x_{1}|z_{1})$
3:for $t=2$ to $T$ do
4:  Receive observation $z_{t}$
5:  $\tilde{x}_{t-1}=\alpha(\tau_{\text{renoise}})\cdot x_{t-1}+\sigma(\tau_{\text{renoise}})\cdot\mathcal{N}(0,I)$
6:  Solve $dx_{t}(\tau)/d\tau=v_{\theta}(x_{t}(\tau),\tau;z_{\leq t}),x_{t}(1)=\tilde{x}_{t-1}$ and obtain $x_{t}=x_{t}(0)$
7:end for
8:return $x_{1},x_{2},\ldots,x_{T}$
@@ 4 Related Works
Bayesian inference of diffusion models. Bayesian inference with diffusion priors has been widely studied, with early work motivated by inverse problems such as image inpainting (Cardoso et al., 2024; Chung et al., 2023; Dou and Song, 2024). The problem is to sample from the posterior $p(x|y)$ given measurements $y=f(x)$ and a diffusion model $p(x)$. See Daras et al. (2024) for a comprehensive review. A natural question is whether Bayesian filtering update (Eqs. 2-3) can be implemented directly for flow-based models based on these approaches. On one hand, these approaches mainly study static inference, while our work focuses on sequential inference in streaming settings under latency constraints. Extending these approaches to streaming settings remains relatively underexplored. Crucially, most existing approaches rely on likelihood-based guidance, requiring explicit, often linear, observation models and system dynamics that are often unavailable in the real world.
@@ 5 Experiment
5.1 Physical System Forecasting
RMSE v.s. Energy Score. We adopt root mean squared error (RMSE) and energy score Gneiting and Raftery (2007) as evaluation metrics, both of which are lower the better. RMSE is a point-wise metric that quantifies prediction error, and the energy score is a distributional metric widely used in probabilistic time-series forecasting to evaluate predictive distributions, defined by $\mathbb{E}_{\hat{x}_{t}\sim p_{\theta}}\norm{\hat{x}_{t}-x_{t}}_{1}-\frac{1}{2}\mathbb{E}_{\hat{x}_{t},\hat{x}_{t}^{\prime}\sim p_{\theta}}\norm{\hat{x}_{t}-\hat{x}_{t}^{\prime}}_{1}$. Basically, the energy score measure the absolute error but with a penalty of over-deterministic prediction. It is known that the expected energy score is minimized if and only if the model prediction exactly recovers the ground-truth distribution (Gneiting and Raftery, 2007). Together, RMSE and the energy score offer complementary perspectives on model performance.
Table 2 reports the RMSE and energy scores of different methods. We observe that a deterministic auto-regressive model (using the same backbone as others) attains RMSE comparable to or better than diffusion and flow matching models. However, it fails to capture the underlying predictive uncertainty, as evidenced by substantially worse energy scores. Asynchronous denoising approaches such as CL-Diffusion intentionally maintain partially denoised future trajectories. For evaluation, we fully denoise these trajectories into clean samples using one-step DDIM. Despite this, their performance remains inferior. Warm-start heuristic exhibits similar behavior, as the generated samples are not guaranteed to align with the target predictive distribution. In contrast, sequential flow models attain both strong point-wise accuracy and well-calibrated predictive distributions using one sampling step.
Table 2: Results of physical system forecasting. Prediction horizon $H=10$ for Burgers’equation and $H=12$ for weather forecasting.
@@ 5.1 Physical System Forecasting
Long-horizon behavior. In appendix C we study the RMSE of different models as a function of the forecast lead time h=1,2,,Hh=1,2,...,H. We find the sequential flow model with one sampling step performs competitively to the full-step pretrained model across the entire forecast window.

### v2

@@ Abstract
Sequential prediction from streaming observations is a fundamental problem in stochastic dynamical systems, where inherent uncertainty often leads to multiple plausible futures. While diffusion and flow-matching models are capable of modeling complex, multi-modal trajectories, their deployment in real-time streaming environments typically relies on repeated sampling from a non-informative initial distribution, incurring substantial inference latency and potential system backlogs. In this work, we introduce Sequential Flow Matching, a principled framework grounded in Bayesian filtering. By treating streaming inference as learning a probability flow that transports the predictive distribution from one time step to the next, our approach naturally aligns with the recursive structure of Bayesian belief updates. We provide theoretical justification that initializing generation from the previous posterior offers a principled warm start that can accelerate sampling compared to naïve re-sampling. Across a wide range of forecasting, decision-making and state estimation tasks, our method achieves performance competitive with full-step diffusion while requiring only one or very few sampling steps, therefore with faster sampling. It suggests that framing sequential inference via Bayesian filtering provides a new and principled perspective towards efficient real-time deployment of flow-based models. Our code is available at https://github.com/Graph-COM/Sequential_Flow_Matching.
Machine Learning, ICML
@@ 1 Introduction
To learn the belief-to-belief probability flow, we propose a flow matching objective function that can be used to effectively finetune a pretrained flow-based model. We evaluate the proposed approach on a diverse set of streaming tasks, including forecasting, planning and control, and state estimation. Our results demonstrate that belief-update-based inference achieves a compelling trade-off between fidelity and efficiency: competitive performance to full-step diffusion with only one or a few sampling steps.
2 Preliminary
@@ 3.1 Problem Setup and Bayesian Filtering
Equations (2),(3) highlight a key structural property of sequential inference: the current belief $p(x_{t}|z_{\leq t})$ can be obtained by propagating the previous belief $p(x_{t-1}|z_{\leq t-1})$ via the dynamics $p(x_{t}|x_{t-1},z_{\leq t-1})$ and a correction using new observation $z_{t}$. Although this exact formulation is often intractable due to unknown system dynamics, Bayesian filtering can be interpreted abstractly as an operator:
@@ 3.2 Sequential Inference with Flow-based Models: Existing Practices
$\frac{d}{d\tau}x_{t}(\tau)=v(x_{t}(\tau),\tau;z_{\leq t}),$ (5)
$x_{t}(0)\sim p(x_{t}|z_{\leq t}),\quad x_{t}(1)\sim\mathcal{N}(0,I),$
@@
To deploy such a model in sequential inference, a common practice is to repeatedly apply the conditional prediction model over time $t$. This receding-horizon practice is widely adopted in time-series forecasting (Box et al., 2015) and decision-making systems such as model predictive control (Garcia et al., 1989). For flow-based models, this naïve practice requires restarting a fresh sampling process and solving ODE Eq. (5) from scratch to incorporate any new observation (Janner et al., 2022).
Consequently, this pipeline suffers from high inference latency for real-time systems: solving ODE Eq. (5) from the non-informative Gaussian demands tens or hundreds of numerical solver iterations and often fails to match the observation arrival frequency, producing potential delayed predictions. This problem becomes more severe when it comes to long-horizon prediction, as each round of estimation is more expensive to compute. Fundamentally, the restarting-sampling paradigm discards historical estimations and, by initializing each update from an independent Gaussian prior, treats temporally correlated physical processes as independent across time.
@@ 3.3 Sequential Flow Matching
From the perspective of Bayesian filtering (Eqs. (2)-(3)), inference over future states should proceed recursively: the belief at time $t$ is obtained by refining the previous belief using new observations. This structure suggests that historical estimations should serve as prior information rather than being discarded. For example, consider a forecasting task where $x_{t}=s_{t+1:t+H}$. The prediction $x_{t-1}=s_{t:t+H-1}$ at time t1t-1 already provides a coarse estimate of future states and naturally approximates $x_{t}=s_{t:t+H}$. Ignoring such information results in “redundant” computation and unnecessary sampling overhead.
Motivated by the recursive structure, we propose to directly parameterize and learn the filtering operator in Eq. (4) via a probability flow. We introduce Sequential Flow Matching model, which learns a probability flow of predictive distribution from time t1t-1 to time $t$, conditioned on newly arrived observations $z_{<t}\rightarrow z_{\leq t}$. The key step is to replace the non-informative Gaussian distribution with the previous estimation $p(x_{t-1}|z_{\leq t-1})$ as the source distribution in flow ODE:
$\frac{d}{d\tau}x_{t}(\tau)=v(x_{t}(\tau),\tau;z_{\leq t}),$ (6)
$x_{t}(0)\sim p(x_{t}|z_{\leq t}),\quad\boldsymbol{x_{t}(1)\sim p(x_{t-1}|z_{\leq t-1})}.$
The process of solving this ODE can be viewed as performing Bayesian filtering in Eq. (4). Figure 1 C illustrates the inference pipeline of sequential flow matching.
@@
(A) independent Gaussian coupling. Suppose $x_{t}(0)\sim p(x_{t}|z_{\leq t})$ and $x_{t}(1)\sim\mathcal{N}(0,I)$ are independently sampled. Let $p_{\text{Gaussian}}$ be the distribution of $\hat{x}_{t}(0)=x_{t}(1)-v(x_{t}(1),1;z_{\leq t})$ using one-step sampling by ODE Eq. (5). Then, $\mathcal{W}^{2}_{2}(p_{\text{Gaussian}},p)=\operatorname{Var}(x_{t}|z_{\leq t}),$ where $\mathcal{W}_{2}(p,q)$ is the 2-Wasserstein distance.
(B) temporally-correlated coupling. Suppose $(x_{t-1},x_{t})\sim p(x_{t-1}|z_{\leq t-1})p(x_{t}|x_{t-1},z_{\leq t})$, and $x_{t}(0)=x_{t},x_{t}(1)=x_{t-1}$. Let $p_{\text{Bayes}}$ be the distribution of one-step sampling by ODE Eq. (6). Then
@@ 3.4 Practical Implementation
In order to learn sequential flow matching (Eq. (6)), the key ingredient is to draw samples of $x_{t-1},x_{t}$ and construct a training dataset. In this section, we discuss practical considerations for dataset construction to train a parameterized velocity $v_{\theta}$ for sequential flow matching.
@@
Distribution alignment via pretrained model rollouts. To reduce train-test mismatch, we leverage a pretrained flow-based model which better approximates $x_{t-1}$ encountered at test time. Concretely, we first pretrain a flow matching $v_{\theta_{0}}$ to learn $p_{\theta_{0}}(x_{t}|z_{\leq t})\approx p(x_{t}|z_{\leq t})$ in a standard sampling-from-Gaussian manner (Eq. (5)). By letting the neural architecture of $v_{\theta_{0}}$ be the same as $v_{\theta}$, we can expect $p_{\theta_{0}}(x_{t-1}|z_{\leq t-1})$ better capture the test-time source distribution $p_{\theta}(x_{t-1}|z_{\leq t-1})$. In practice, we construct pairs $(x_{t-1},x_{t})$ where $x_{t-1}$ comes from pretrained model and $x_{t}$ comes from ground truth. This construction aligns the source distribution with the model predictive distribution encountered at test time, while avoiding compounding errors that would arise from fully synthetic trajectories. The constructed dataset is used to finetune $\theta$ initialized from $\theta_{0}$. We found this hybrid construction effectively reduce train-test mismatch and adopt it throughout our experiments.
Re-noising mechanism. While using generated data from pretrained model reduces train-test mismatch, the small deviations in model prediction at test-time can still accumulate and lead to compounding errors over multiple rounds of belief update. To mitigate the error accumulation, we adopt a re-noising mechanism that injects noise into the previous estimation of $x_{t-1}$ at both training and test time, before applying the flow. Concretely, in flow ODE Eq. (6), we replace source distribution $p(x_{t-1}|z_{\leq t-1})$ with a noisy version $\tilde{x}_{t-1}:=\alpha(\tau_{\text{renoise}})x_{t-1}+\sigma(\tau_{\text{renoise}})\cdot\mathcal{N}(0,I)$, where $\alpha(\tau_{\text{renoise}}),\sigma(\tau_{\text{renoise}})$ are the coefficients of the probability flow path at noise level (flow time) $\tau_{\text{renoise}}$. We treat $\tau_{\text{renoise}}$ as a hyperparameter that can be tuned in practice.
The re-noise level $\tau_{\text{renoise}}$ controls the trade-offs between preserving historical prediction information for sampling acceleration and accommodating uncertainty during belief refinement. If $\tau_{\text{renoise}}=1$, we have $\tilde{x}_{t-1}\sim\mathcal{N}(0,I)$, which reduces to the restarting-sampling paradigm that requires a large number of sampling steps. If $\tau_{\text{renoise}}=0$, it relies on a precise modeling of $x_{t-1}\sim p(x_{t-1}|z_{\leq t-1})$, which might induces error accumulation for long-horizon prediction. In Appendix C.3 we study the impact of $\tau_{\text{renoise}}$, and we observe that a suitable re-noise level varies with the level of environmental stochasticity.
We present the finetuning algorithm (Algorithm 1) and inference algorithms (Algorithm 2).
@@
0: Ground-truth $x_{t}$, context $z_{\leq t}$, renoise level $\tau_{\text{renoise}}$, and fixed pretrained model $v_{\theta_{0}}$ (from Eq. (5))
1: Generate $x_{t-1}\sim p_{\theta_{0}}(x_{t-1}|z_{\leq t-1})$ using $v_{\theta_{0}}$
2: Renoise $\tilde{x}_{t-1}=\alpha(\tau_{\text{renoise}})\cdot x_{t-1}+\sigma(\tau_{\text{renoise}})\cdot\mathcal{N}(0,I)$
3: Sample interpolation time $\tau\sim\text{Uniform}(0,1)$
4: Compute $x_{t}(\tau)=\alpha(\tau)x_{t}+\sigma(\tau)\tilde{x}_{t-1}$
5:Return loss $L(\theta)=\norm{v_{\theta}(x_{t}(\tau),\tau;z_{\leq t})-\dot{x}_{t}(\tau)}^{2}$
@@
0: Pretrained model $v_{\theta_{0}}$ (from Eq. (5)), sequential flow matching model $v_{\theta}$ finetuned by Algorithm 1, initial observation $z_{1}$
1: initialize $x_{1}\sim p_{\theta_{0}}(x_{1}|z_{1})$
2:for $t=2$ to $T$ do
3:  Receive observation $z_{t}$
4:  $\tilde{x}_{t-1}=\alpha(\tau_{\text{renoise}})\cdot x_{t-1}+\sigma(\tau_{\text{renoise}})\cdot\mathcal{N}(0,I)$
5:  Solve $dx_{t}(\tau)/d\tau=v_{\theta}(x_{t}(\tau),\tau;z_{\leq t}),x_{t}(1)=\tilde{x}_{t-1}$ and obtain $x_{t}=x_{t}(0)$
6:end for
7:Return $x_{1},x_{2},\ldots,x_{T}$
@@ 4 Related Works
Bayesian inference of diffusion models. Bayesian inference with diffusion priors has been widely studied, with early work motivated by inverse problems such as image inpainting (Cardoso et al., 2024; Chung et al., 2023; Dou and Song, 2024). The problem is to sample from the posterior $p(x|y)$ given measurements $y=f(x)$ and a diffusion model $p(x)$. See Daras et al. (2024) for a comprehensive review. A natural question is whether Bayesian filtering update (Eqs. (2)-(3)) can be implemented directly for flow-based models based on these approaches. On one hand, these approaches mainly study static inference, while our work focuses on sequential inference in streaming settings under latency constraints. Extending these approaches to streaming settings remains relatively underexplored. Crucially, most existing approaches rely on likelihood-based guidance, requiring explicit, often linear, observation models and system dynamics that are often unavailable in the real world.
@@ 5 Experiment
Table 2: Results of physical system forecasting. NFE refers to number of function evaluation (sampling steps). Prediction horizon $H=10$ for Burgers’equation and $H=12$ for weather forecasting.
@@
5.1 Physical System Forecasting
RMSE v.s. Energy Score. We adopt root mean squared error (RMSE) and energy score (Gneiting and Raftery, 2007) as evaluation metrics, both of which are lower the better. RMSE is a point-wise metric that quantifies prediction error, and the energy score is a distributional metric widely used in probabilistic time-series forecasting to evaluate predictive distributions, defined by $\mathbb{E}_{\hat{x}_{t}\sim p_{\theta}}\norm{\hat{x}_{t}-x_{t}}_{1}-\frac{1}{2}\mathbb{E}_{\hat{x}_{t},\hat{x}_{t}^{\prime}\sim p_{\theta}}\norm{\hat{x}_{t}-\hat{x}_{t}^{\prime}}_{1}$. Basically, the energy score measure the absolute error but with a penalty of over-deterministic prediction. It is known that the expected energy score is minimized if and only if the model prediction exactly recovers the ground-truth distribution (Gneiting and Raftery, 2007). Together, RMSE and the energy score offer complementary perspectives on model performance.
Table 2 reports the RMSE and energy scores of different methods. We observe that a deterministic auto-regressive model (using the same backbone as others) attains RMSE comparable to or better than diffusion and flow matching models. However, it fails to capture the underlying predictive uncertainty, as evidenced by substantially worse energy scores. Asynchronous denoising approaches such as CL-Diffusion intentionally maintain partially denoised future trajectories. For evaluation, we fully denoise these trajectories into clean samples using one-step DDIM. Despite this, their performance remains inferior. Warm-start heuristic exhibits similar behavior, as the generated samples are not guaranteed to align with the target predictive distribution. In contrast, sequential flow models attain both strong point-wise accuracy and well-calibrated predictive distributions using only one sampling step.
Long-horizon behavior. Figure 2 compares the RMSE as a function of the forecast lead time h=1,2,,Hh=1,2,...,H. While the auto-regressive model performs better at short horizons, its error increases more rapidly with horizon than flow-based models. In contrast, the sequential flow model with one sampling step performs competitively to the full-step pretrained model across the entire forecast window.
Figure 2: Prediction RMSE of Burgers’ equation as a function of forecast lead time (number of time-steps predicted ahead).